OpenStudy (anonymous):

Solve x – 2y = 4 for y.

OpenStudy (anonymous):

subtract \(x\) from both sides divide both sides by \(-2\)

OpenStudy (anonymous):

let me know what you get

OpenStudy (anonymous):

y=-2-x

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

\[x-2y=4\] \[-2y=4-x\] \[y=\frac{4-x}{-2}\]which you can rewrite as \[y=\frac{x-4}{2}\] if you like or even \[y=\frac{1}{2}x-2\]

OpenStudy (anonymous):

but you have to divide the \(x\) as well as the \(4\)

OpenStudy (anonymous):

jeez i really messed that one up lol thankyou!

OpenStudy (anonymous):

yw